# LOD-GS: Level-of-Detail-Sensitive 3D Gaussian Splatting for Detail Conserved Anti-Aliasing

### v2

@@ 4.2 EWA Filtering
In the original implementation of 3DGS, a dilation operator [50] is applied to each 2D splat. This helps avoid small primitives that are hard to optimize. The dilation operator could be written as: $\mathcal{G}_{k}^{2D}(\mathrm{x})=e^{-\frac{1}{2}(\mathrm{x}-\mathbf{p}_{k})^{T}(\Sigma_{k}^{2D}+s\mathbf{I})(\mathrm{x}-\mathbf{p}_{k})}$. This operator leads to obvious dilation artifacts when the image resolution decreases. These artifacts are primarily caused by the expansion of the primitive’s scale without adjusting its opacity. As a result, the overall energy increases, leading to the dilation effect. Mip-Splatting [50] utilizes EWA Filtering [55] to alleviate this artifact:
@@
The main difference between dilation operator and EWA filter is the normalization term preceding the exponential term. In LOD-GS, we also use the EWA filter before the alpha blending computation to reduce the aliasing artifact. After the EWA filtering, we perform alpha blending to obtain the rendering result as follows:
$\mathbf{c}(\mathbf{x})$ $=\sum_{k=1}^{K}\left(\mathbf{c}_{k}+\mathcal{F}_{c}(\nu)\right)\hat{\alpha_{k}}\prod_{j=1}^{k-1}(1-\hat{\alpha_{j}})$ (9)
$\hat{\alpha}_{k}$ $=\sqrt{\frac{|\Sigma_{k}^{2D}|}{|\Sigma_{k}^{2D}+s\mathbf{I}|}}(\alpha_{k}+\mathcal{F}_{\alpha}(\nu))$ (10)
where $\hat{\alpha}$ is the opacity of the filtered Gaussian primitive, and $\mathcal{F}_{c}$ is a learnable module that models the slight color differences caused by changes in the sampling rate. After obtaining the rendering results, we compute the image loss between these results and the ground truth. Then we perform gradient descent to optimize all Gaussian parameters and the filter prediction module in an end-to-end manner.
Figure 4: Qualitative comparison on our extended Blender dataset. All methods are trained at three levels (near, middle, and far) and tested at the near level. Our method more effectively captures the details of microstructures and textures.
@@
Level 1 Level 2 Level 3 Avg. Level 1 Level 2 Level 3 Avg. Level 1 Level 2 Level 3 Avg.
3DGS [14] 27.20 32.57 37.12 32.30 0.906 0.970 0.995 0.957 0.150 0.031 0.005 0.062
Mipmap-GS [19] 26.59 31.61 34.66 30.95 0.898 0.967 0.992 0.952 0.157 0.034 0.005 0.066
Multiscale-3DGS[45] 26.31 31.54 34.66 30.84 0.898 0.967 0.992 0.952 0.158 0.034 0.005 0.066
Mip-Splatting [50] 26.87 33.08 40.76 33.57 0.902 0.973 0.997 0.957 0.154 0.029 0.003 0.062
Analytic-Splatting [21] 27.06 33.27 40.41 33.58 0.901 0.972 0.997 0.957 0.154 0.030 0.003 0.062
@@ 5.1 Implementation
The implementation of LOD-GS follows the framework of vanilla 3DGS, which is implemented using PyTorch and CUDA code. We developed a customized CUDA rasterizer to support the forward and backward of our LOD-GS. The primitive level appearance change is fitted using a group of Gaussian functions, we set the number of these Gaussians to 20 in the synthetic dataset, and set it to 8 in Mip-NeRF 360 dataset for a smaller model size so that our machine can perform training. We evaluate our method on public datasets, including the Multi-scale Synthetic Dataset from Mip-NeRF [2] and the Mip-NeRF 360 [3] dataset. This allows us to test our performance on both synthetic scenes and real-world scenes. We also render a multi-level synthetic dataset for evaluation as illustrated in Section 5.3. All experiments are conducted on a single RTX3090 GPU.
@@ 5.2 Evaluation on the Multi-scale Blender Dataset
The Blender dataset introduced in the original NeRF [27] is a synthetic dataset where all training and testing images observe the scene content from a roughly constant distance with the same focal length, which differs significantly from real-world captures. MipNeRF [1] introduces a multi-scale Blender dataset designed to enhance the evaluation of reconstruction accuracy and anti-aliasing in multi-resolution scenes. This dataset is generated by downscaling the original dataset by factors of 2, 4, and 8, and then combining these variations. We evaluate our LOD-GS on this dataset with several competitive methods, including NeRF-based methods (NeRF [27], MipNeRF [1], and Tri-MipRF [12]) and 3DGS-based methods (Mipmap-GS [18], Multiscale-3DGS [45], Mip-Splatting [50], and Analytic-Splatting [21]). Following previous works, we report three metrics: PSNR, SSIM [39], and VGG LPIPS [52] across four resolutions and average results, as shown in Table 3.2. Mip-NeRF and Tri-MipRF exhibit stronger anti-aliasing capabilities compared to vanilla NeRF. For 3DGS-based methods, Mip-Splatting and Analytic-Splatting significantly enhance the anti-aliasing ability of 3DGS. While Multiscale-3DGS can achieve better performance at the original resolution, it degenerates at lower resolutions. It is important to emphasize that both Mip-Splatting and Analytic-Splatting require sampling more cases from higher resolutions to achieve good performance. In contrast, by utilizing LOD-sensitive filter control, our method achieves rather good performance without relying on any dataset sampling technique. As shown in Table 3.2, our LOD-GS achieves state-of-the-art anti-aliasing performance while maintaining competitive rendering quality at the original resolution. To test the effectiveness of the modules in LOD-GS, we also performed an ablation study. The experimental results are summarized in the last two rows of the Table 3.2. As shown, removing either the EWA filter or the LOD filter leads to a performance drop, especially at lower resolutions. This indicate that both these two modules are crucial for the anti-aliasing of 3DGS.
@@ 5.3 Evaluation on the Multi-level Blender Dataset
Both decreasing the focal length and increasing the camera distance can change the sampling rate, thereby reducing the resolution of the photographed object. Mip-NeRF reduces image resolution using image processing techniques, attributing this change to alterations in focal length. However, in 3D applications, the most common approach to change the sampling rate is by adjusting the camera distance, rather than the focal length. There is a difference between adjusting the camera distance and the focal length. To make a more comprehensive evaluation, we re-render the synthetic dataset used in NeRF [27] from three different camera distances. We call this extended dataset the “Multi-level Blender Dataset” and assign Level 1 to the camera with the nearest camera distance to the object, Level 2 to the middle camera distance, and Level 3 to the farthest distance. The results of the comparative experiments on this dataset are presented in Table 4.2.
@@
In addition to capturing fine details, we also evaluate the anti-aliasing capabilities of our method. We render the object from a considerable distance to assess this ability. The qualitative results are illustrated in Figure 5. As shown, 3DGS tends to generate dilated boundaries and aliased results in areas with high-frequency texture. In contrast, both Mip-Splatting and our LOD-GS effectively handle aliasing when the camera is far from the scene, i.e., at a low sampling rate. Compared to Mip-Splatting, our method can reconstruct more details, such as the inner structure of the microphone, as shown in the third row of Figure 5.
@@ 5.4 Evaluation on the Mip-NeRF 360 Dataset
To test the performance of our method on real-world data, we evaluate our method and perform comparisons on the Mip-NeRF 360 Dataset [3]. Due to our limited computational resources, we cannot perform training on all resolutions in the dataset. Therefore, we preprocess the full-resolution images into three downsampled scales ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$) and conduct multi-scale training and testing. The qualitative and quantitative experimental results are shown in Figure 6 and Table 5.4, respectively. As shown in Table 5.4, our method achieves SOTA performance across nearly all metrics at different resolutions.
Figure 6: Qualitative comparison on the Mip-NeRF 360 dataset [3]. All methods are trained and tested at three downsampled resolutions ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$). We select two scenes (garden and bicycle) to compare the high and low-resolution rendering results.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$\nicefrac{{1}}{{8}}$ Res. $\nicefrac{{1}}{{16}}$ Res. $\nicefrac{{1}}{{32}}$ Res. Avg. $\nicefrac{{1}}{{8}}$ Res. $\nicefrac{{1}}{{16}}$ Res. $\nicefrac{{1}}{{32}}$ Res. Avg. $\nicefrac{{1}}{{8}}$ Res. $\nicefrac{{1}}{{16}}$ Res. $\nicefrac{{1}}{{32}}$ Res. Avg.
3DGS [14] 27.99 29.96 29.62 29.19 0.850 0.919 0.934 0.901 0.158 0.074 0.057 0.096
Mipmap-GS [18] 27.28 29.23 28.50 28.33 0.836 0.909 0.919 0.888 0.173 0.084 0.066 0.108
Multiscale-3DGS[45] 27.30 29.25 28.50 28.35 0.836 0.910 0.920 0.888 0.173 0.084 0.066 0.108
Mip-Splatting [50] 28.35 30.16 31.06 29.86 0.865 0.924 0.949 0.913 0.132 0.064 0.043 0.080
Analytic-Splatting [21] 28.84 30.63 31.63 30.37 0.868 0.926 0.953 0.915 0.128 0.066 0.043 0.079
LOD-GS (ours) 28.96 30.90 32.40 30.75 0.869 0.929 0.958 0.919 0.134 0.065 0.038
Table 3: Multi-scale training and Multi-scale testing on the Mip-NeRF 360 dataset [3]. All methods are trained and tested at three downsampled resolutions ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$). Our model achieves state-of-the-art performance across most metrics in comparison.
Qualitative Results    To provide a more intuitive comparison on real-world datasets, we analyze the rendering results from different resolutions and methods. As shown in Figure 6, the $\nicefrac{{1}}{{8}}$ garden images reveal that both 3DGS and Multiscale-3DGS exhibit a blurred appearance, lacking fine details. While Mip-Splatting captures some details, its rendering results are overly smooth. Analytic-Splatting effectively learns the detailed textures of objects, but our method achieves higher rendering quality than Analytic-Splatting. It is important to note that the additional workload introduced by Analytic-Splatting limits its rendering speed to around 30 FPS, whereas our method performs efficiently at approximately 45 FPS. In addition to its ability to reconstruct fine details, our LOD-GS also possesses strong anti-aliasing capabilities. Examining the rendering results of bicycle spokes at the $\nicefrac{{1}}{{32}}$ downsampled resolution, we see that 3DGS generates aliased results. MSGS reduces aliasing to some degree but introduces dilation artifacts. Mip-Splatting and Analytic-Splatting manage to remove aliasing but sacrifice the geometric details of the bicycle spokes. In contrast, LOD-GS generates better anti-aliased results while maximizing the preservation of the original image structure. Based on both quantitative and qualitative comparisons, our proposed LOD-GS achieves detail-preserved and aliasing-free rendering at the same time.
6 Conclusion
In this paper, we propose a Level-of-Detail-Sensitive Gaussian Splatting method, LOD-GS, to sense and learn the appearance changes caused by variations in the camera sampling rate. Compared to existing methods, our approach achieves state-of-the-art performance on anti-aliasing tasks while preserving fine details. It alleviates the ambiguity in multi-scale training by disentangling inputs with different sampling rates. To perform a comprehensive evaluation, we re-render the Synthetic NeRF dataset from different camera distances to account for sampling rate changes caused by camera distance. Experimental results demonstrate that LOD-GS can utilize multi-scale 2D images to train a prefiltered radiance field, allowing the model to display detailed textures without aliasing across different sampling rates. LOD-GS is expected to further improve the reconstruction quality of neural radiance fields, especially when the collected images are captured at varying sampling rates.

### v3

@@ 4.2 EWA Filtering
In the implementation of 3DGS, a dilation operator [mip-splatting] is applied to each projected Gaussian. This helps avoid small primitives that are hard to optimize. The dilation operator could be written as: $\mathcal{G}_{k}^{2D}(\mathrm{x})=e^{-\frac{1}{2}(\mathrm{x}-\mathbf{p}_{k})^{T}(\Sigma_{k}^{2D}+s\mathbf{I})(\mathrm{x}-\mathbf{p}_{k})}$. This operator leads to obvious dilation artifacts when the image resolution decreases. These artifacts are primarily caused by the expansion of the primitive’s scale without adjusting its opacity. As a result, the overall energy increases, leading to the dilation effect. The EWA Filtering [zwicker2001ewa] could be used to alleviate this artifact:
@@
The main difference between dilation operator and EWA filter is the normalization term preceding the exponential term. In LOD-GS, we use the EWA filter before the alpha blending computation to reduce the dilation and aliasing artifact. After the EWA filtering, we perform alpha blending to obtain the rendering result as follows:
$\mathbf{c}(\mathbf{x})$ $=\sum_{k=1}^{K}\mathbf{c}_{k}\hat{\alpha_{k}}\mathcal{G}_{k}^{2D}(\mathrm{x})\prod_{j=1}^{k-1}(1-\hat{\alpha_{j}}\mathcal{G}_{j}^{2D}(\mathrm{x}))$ (9)
$\hat{\alpha}_{k}$ $=\alpha_{k}+\mathcal{F}_{\alpha}(\nu)$ (10)
where $\hat{\alpha}$ is the opacity of the LOD-filtered Gaussian primitive. After obtaining the rendering results, we compute the image loss between these results and the ground truth. Then we perform gradient descent to optimize all Gaussian parameters and the filter prediction module in an end-to-end manner. This design enables the LOD filter to adjust its filtering degree by taking the EWA filter into account, allowing it to predict the most appropriate filter for achieving both aliasing-free and detail-conserved rendering. Conversely, the 3D filter and 2D Mip filter of Mip-Splatting operate sequentially but independently, which can easily result in over-filtering and lead to blurry rendering results.
Figure 4: Qualitative comparison on our extended Blender dataset. All methods are trained across three levels: $L_{1}$ (near), $L_{2}$ (middle), and $L_{3}$ (far). Our method more effectively captures the details of microstructures and textures during the multi-level training.
@@
Near Medium Far Avg. Near Medium Far Avg. Near Medium Far Avg.
3DGS [kerbl3Dgaussians] 27.20 32.57 37.12 32.30 0.906 0.970 0.995 0.957 0.150 0.031 0.005 0.062
Mipmap-GS [mipmapgs] 26.59 31.61 34.66 30.95 0.898 0.967 0.992 0.952 0.157 0.034 0.005 0.066
Multiscale-3DGS[multiscale-3dgs] 26.31 31.54 34.66 30.84 0.898 0.967 0.992 0.952 0.158 0.034 0.005 0.066
Mip-Splatting [mip-splatting] 26.87 33.08 40.76 33.57 0.902 0.973 0.997 0.957 0.154 0.029 0.003 0.062
Analytic-Splatting [analytic-splatting] 27.06 33.27 40.41 33.58 0.901 0.972 0.997 0.957 0.154 0.030 0.003 0.062
@@ 5.1 Implementation
The LOD filtering is implemented using PyTorch and is designed to be plug-and-play, allowing for easy integration into existing 3DGS-based methods. It is implemented using GMM and we set the number of basis functions to 20 in the following experiment. We evaluate our method on both synthetic and real-world datasets, including the Multi-scale Synthetic Dataset from Mip-NeRF [mip-nerf] and the Mip-NeRF 360 [mipnerf360] dataset. We render a multi-level synthetic dataset for more comprehensive evaluation as illustrated in Section 5.3. We also evaluate the generalization ability of our method in Section LABEL:subsec:stmt and examine the impact of the number of basis functions used in the LOD filter in Section LABEL:subsec:basis. In our experiment, Mip-Splatting uses the densification scheme from GOF [GOF], while all other methods, including LOD-GS, utilize the original densification scheme of 3DGS. All experiments are conducted on a single RTX3090 GPU.
@@ 5.2 Evaluation on the Multi-scale Blender Dataset
The Blender dataset introduced in the original NeRF [mildenhall2020nerf] is a synthetic dataset where all training and testing images observe the scene content from a roughly constant distance with the same focal length, which differs significantly from real-world captures. MipNeRF [barron2021mip] introduces a multi-scale Blender dataset designed to enhance the evaluation of reconstruction accuracy and anti-aliasing in multi-resolution scenes. This dataset is generated by downscaling the original dataset by factors of 2, 4, and 8, and then combining these variations. We evaluate our LOD-GS on this dataset with several competitive methods, including NeRF-based methods (NeRF [mildenhall2020nerf], MipNeRF [barron2021mip], and Tri-MipRF [trimip-nerf]) and 3DGS-based methods (Mipmap-GS [mipmapgs], SA-GS [sa-gs], Multiscale-3DGS [multiscale-3dgs], Mip-Splatting [mip-splatting], and Analytic-Splatting [analytic-splatting]). Following previous works, we report three metrics: PSNR, SSIM [wang2004image], and VGG LPIPS [zhang2018unreasonable] across four resolutions and average results, as shown in Table 3.2. Mip-NeRF and Tri-MipRF exhibit stronger anti-aliasing capabilities compared to vanilla NeRF. For 3DGS-based methods, Mip-Splatting and Analytic-Splatting significantly enhance the anti-aliasing ability of 3DGS. While Multiscale-3DGS can achieve better performance at the original resolution, it degenerates at lower resolutions. It is important to emphasize that both Mip-Splatting and Analytic-Splatting require sampling more cases from high resolutions to maintain their performance at the original resolution. In contrast, by utilizing LOD-sensitive filter control, our method achieves rather good performance without relying on any dataset resampling trick. As shown in Table 3.2, our LOD-GS achieves state-of-the-art anti-aliasing performance while maintaining competitive rendering quality at the original resolution.
@@ 5.3 Evaluation on the Multi-level Blender Dataset
Both decreasing the focal length and increasing the camera distance can change the sampling rate, thereby reducing the resolution of the photographed object. Mip-NeRF reduces image resolution using image processing techniques, attributing this change to alterations in focal length. However, in 3D applications, the most common approach to change the sampling rate is by adjusting the camera distance, rather than the focal length. There is a difference between adjusting the camera distance and the focal length. To make a more comprehensive evaluation, we re-render the synthetic dataset used in NeRF [mildenhall2020nerf] from three different camera distances. We call this extended dataset the “Multi-level Blender Dataset” and assign Level 1 to the camera with the nearest camera distance to the object, Level 2 to the middle camera distance, and Level 3 to the farthest distance. The results of the comparative experiments on this dataset are presented in Table 4.2. As shown, 3DGS performs well when the object is close to the camera but degrades in performance when the camera is far from the object. Conversely, Mip-Splatting and Analytic-Splatting exhibit better performance at greater distances but struggle at near distances. Only LOD-GS achieves good performance at both near and far distances. It is crucial to note that when the camera is positioned at a far distance, a significant portion of the image appears white. This can lead to smaller differences in SSIM and LPIPS between various methods. In this context, PSNR is a more appropriate metric, as it is calculated based on pixel differences.
@@
In addition to capturing fine details, we also evaluate the anti-aliasing capabilities of our method. We render the object from a considerable distance to assess this ability. The qualitative results are illustrated in Figure 5. As shown, 3DGS tends to generate dilated boundaries and aliased results in areas with high-frequency texture. In contrast, both Mip-Splatting and our LOD-GS effectively handle aliasing when the camera is far from the scene, i.e., at a low sampling rate. Compared to Mip-Splatting, our method can reconstruct more details, such as the inner structure of the microphone and the rope net of the ship, as shown in the third row and the fourth row of the Figure 5.
@@ 5.4 Evaluation on the Mip-NeRF 360 Dataset
To test the performance of our method on real-world data, we evaluate our method and perform comparisons on the Mip-NeRF 360 Dataset [mipnerf360]. Due to the high resolution of images in the 360 Dataset and the need to perform multiscale training, we conduct multiscale training and testing on three downsampled scales ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$) to reduce the VRAM consumption. The qualitative and quantitative experimental results are shown in Figure 6 and Table LABEL:tab:multi_360. As shown in Table LABEL:tab:multi_360, our method achieves SOTA performance across nearly all metrics at different resolutions.
Figure 6: Qualitative comparison on the Mip-NeRF 360 dataset [mipnerf360]. All methods are trained and tested at three downsampled resolutions ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$). We recommend that readers scale up this image and compare the region marked with red boxes across different methods.
Detail Conserved Anti-Aliasing
Supplementary Material
In this supplementary material, we provide a comprehensive ablation study and its analysis in Section 7. We also report the complete results of comparison experiments conducted on the Multi-scale Blender Dataset [mip-nerf], our newly rendered Multi-level Blender Dataset, and the Mip-NeRF 360 Dataset [mipnerf360] in Section 8.
7 Ablation
The capability of our method to achieve aliasing-free and detail-conserved rendering relies on two components: the level-of-detail sensitive filter (denoted as LOD filter hereafter) and the EWA filter. To assess the contribution of these modules to overall performance of our method, we separately remove the LOD and EWA filters and conduct model training on synthetic and real-world dataset. Detailed analyses of these two components are presented in Section 7.1 and Section 7.2.
7.1 Effectiveness of the LOD filter
The LOD filter is mainly responsible for disentangling inputs with different sampling rates. It helps alleviate the ambiguity caused by the multi-scale input, which is the key to learn a radiance field with fine details and anti-aliasing ability at the same time. As shown in Figure 8, scenes trained without the LOD filter fail to capture the fine details of the objects. For instance, the inner structure of the microphone is unclear when photographed from a near distance. Other details, such as the metal accents and rope net of the ship, as well as the bumps on the chair, cannot be well reconstructed without the LOD filter. Moreover, as illustrated in Figure 9, using only the LOD filter can mitigate aliasing to some degree compared to the vanilla 3DGS [kerbl3Dgaussians]. Experimental results in Table 3.2 and Table 3.2 demonstrate that removing the LOD filter leads to a more significant performance drop compared to removing the EWA filter. All these results illustrate that the LOD filter in our method is crucial for anti-aliasing and the conservation of fine details.
7.2 Effectiveness of the EWA filter
The EWA filter is utilized to enhance the model’s anti-aliasing capability. Although our LOD filter can address the aliasing problem to some extent, integrating the EWA filter into our method can provide improved anti-aliasing performance. As shown in Figure 8, the results rendered from a far distance using the method that only utilizes the LOD filter face issues of aliasing and dilation artifacts. More specifically, the result for mic is aliased, while the results for ship and lego exhibit dilation artifacts. The quantitative results in Table 3.2 and Table 3.2 illustrate that removing the EWA filter also leads to a performance drop. These observations and results demonstrate that the EWA filter contributes to the anti-aliasing capability of our method.
8 Additional Results
In this section, we report the detail of our comparison experiment on each scene across different sampling rates. We analyze the experimental results on the Blender Dataset in Section 8.1 and the Mip-NeRF 360 Dataset in Section 8.2.
8.1 Blender Dataset
We perform comparison experiments on the Multi-Scale Blender Dataset [mip-nerf] and our newly rendered Multi-Level Dataset. The Multi-Level Blender dataset does not contain the drum scene because of the wrong specular reflection effect in the drums blender project. The experimental results on the Multi-Scale Blender dataset are shown in Table 3.2. Compared to existing powerful methods for radiance field anti-aliasing, such as Mip-NeRF [mip-nerf], Mip-Splatting [mip-splatting], and Analytic-Splatting [analytic-splatting], our method achieves state-of-the-art performance in most scenes. We collected a new dataset called the Multi-Level Blender dataset, which simulates changes in sampling rate using varying camera distances. The experimental results on this new dataset are reported in Table 3.2. We compare our method with Gaussian Splatting-based methods designed to handle multiscale inputs, including Mipmap-GS [mipmapgs], MSGS [multiscale-3dgs], Mip-Splatting [mip-splatting], and Analytic-Splatting [analytic-splatting]. As shown in Table 3.2, our method achieves the best performance in most scenes. These experimental results demonstrate the state-of-the-art performance of our method in handling inputs with varying sampling rates.
8.2 Mip-NeRF 360 Dataset
We down-sample the full resolution images in Mip-NeRF 360 dataset in three scales ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{32}}$). All methods are trained and tested on these three scales. The experimental results are displayed in the Table 3.2. As shown, our proposed LOD-GS achieves the best or nearly the best performance across all scenes. This illustrates that our method is also applicable to wild and unbounded real-world data.
Figure 8: Qualitative Results of the Ablation Study on the Blender Dataset [mildenhall2020nerf, mip-nerf]. We render the objects from both far and near distances to test the anti-aliasing and detail conservation abilities.
Figure 9: Qualitative Results of the Ablation Study on the Mip-NeRF 360 Dataset [mipnerf360]. All methods are trained and tested on multiscale inputs ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, $\nicefrac{{1}}{{32}}$). We present rendering results from different methods at $\nicefrac{{1}}{{8}}$ and $\nicefrac{{1}}{{32}}$ scales to evaluate detail conservation and anti-aliasing performance.
Table 6: Quantitative Results of the Ablation Study on the Blender Dataset [mildenhall2020nerf, mip-nerf]. We present the experimental results of our full model alongside methods that exclude the LOD module or EWA filter across different scenes and resolutions. The results demonstrate that both the LOD and EWA filters significantly contribute to the overall performance of our method.
Table 7: Quantitative Results of the Ablation Study on the Mip-NeRF 360 Dataset [mipnerf360]. All methods are trained and tested on multiscale inputs. We present the experimental results of our full model alongside methods that exclude either the LOD module or the EWA filter across different scenes and resolutions.
Table 8: Multi-scale Training and Multi-scale Testing on the the Blender dataset [mildenhall2020nerf]. For each scene, we report the arithmetic mean of each metric averaged over the four scales used in the dataset (full resolution, $\nicefrac{{1}}{{2}}$, $\nicefrac{{1}}{{4}}$, and $\nicefrac{{1}}{{8}}$ downsampled scales).
Table 9: Multi-Level Training and Multi-Level Testing on our extended Blender dataset [mildenhall2020nerf]. For each scene, we report the arithmetic mean of each metric averaged over the 3 Levels used in the dataset (near, middle and far).
Table 10: Multi-Scale Training and Multi-Scale Testing on the the Mip-NeRF 360 dataset [mipnerf360]. For each scene, we report the arithmetic mean of each metric averaged over the 3 scales used in the dataset ($\nicefrac{{1}}{{8}}$, $\nicefrac{{1}}{{16}}$, and $\nicefrac{{1}}{{32}}$ downsampled scales).